OpenStudy (anonymous):

how to simplify 5(x-2)-(x+6)+3x can you guys help pleasee!!!!!! (show work so i can know )😁😘

OpenStudy (steve816):

First things first. In order to simplify, you first need to distribute.

OpenStudy (steve816):

5(x-2) = 5x-19 -(x+6) = -x - 6 So we get 5x-10-x-6+3x

OpenStudy (steve816):

Now we just add or subtract like terms! The answer we get is 7x-16 There you go! That is the simplified answer :)

OpenStudy (anonymous):

thank youu😘

OpenStudy (steve816):

No problem my friend.